OpenStudy (anonymous):

WHY WONT ANYONE HELP ME?! how tall is a building that casts a 69 ft shadow when the angle of the suns elevation is 29 degrees

OpenStudy (ilovecake):

I will help you

OpenStudy (anonymous):

me too :D

OpenStudy (jadeishere):

this is a Pythagorean theorem problem :D

OpenStudy (ilovecake):

ok so first......

OpenStudy (ilovecake):

Let me find a website to help you

OpenStudy (anonymous):

i just need a awnser.

OpenStudy (anonymous):

you shouldn't ask for just answers, you wouldn't learn that way :(

OpenStudy (anonymous):

besides, its against the code of conduct :( http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

and?

OpenStudy (anonymous):

i just need an awnser for only one and the next one i ask you guys iill ask fior explanation

OpenStudy (ilovecake):

there

OpenStudy (anonymous):

there you go... the answer's on there, but I recommend you actually read the explanation. @Ilovecake thank u :D

OpenStudy (ilovecake):

Your welcome:)

OpenStudy (anonymous):

:D it is a very good explanation

OpenStudy (anonymous):

thats not the awnser to mine

OpenStudy (anonymous):

it shows you how to do it...

OpenStudy (anonymous):

awnser to mine?

OpenStudy (ilovecake):

Its not suppoesed to give you the EXACT answer you just need to read it and understand it

OpenStudy (anonymous):

I JUST NEED A AWNSER

OpenStudy (ilovecake):

wow

OpenStudy (ilovecake):

I am just trying to help

OpenStudy (anonymous):

THANKS BUT I NEED ONLY A AWNSER

OpenStudy (anonymous):

Tangent (theta) = opposite / adjacent Tangent (69) = X / 69 X = ? I set it up for you, try to figure out the rest yourself. you shouldn't get just answers... you will never learn.

OpenStudy (anonymous):

i apriteate your help thank you for your time

OpenStudy (ilovecake):

I am not going to GIVE YOU ANSWERS

OpenStudy (anonymous):

Tangent (29) = X / 69... actually

OpenStudy (anonymous):

then leave!

OpenStudy (anonymous):

i got a weird desimal

OpenStudy (anonymous):

you should be polite to those trying to teach you. Ilovecake was helping you... and what did you get? let me take a look at it...

OpenStudy (ilovecake):

you have some nerves. No ONE IS GOING TO GIVE YOU ANSWERS JUST BECAUSE YOU WANT SOMEONE TO GIVE IT TO YOU. Just close the question no one is going to help you like that.

OpenStudy (anonymous):

0.420

OpenStudy (anonymous):

29/69

OpenStudy (ilovecake):

@AccessDenied

OpenStudy (ilovecake):

@undeadknight26

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@thomaster

undeadknight26 (undeadknight26):

Sorry i don't give answers.

OpenStudy (ilovecake):

He WANTED the answers. I gave him a website and he said I only need answers. can you give him an explanation

OpenStudy (ilovecake):

He closed the question so, I can leave

OpenStudy (anonymous):

yes you can. thank you for being a good user of OS :D

OpenStudy (ilovecake):

:)